OpenStudy (anonymous):

Instructions:Type the correct answer in the box. Use numerals instead of words. If necessary, use / for the fraction bar. This data set represents the number of cups of coffee sold in a café between 8 a.m. and 10 a.m. every day for 14 days. {10, 9, 6, 12, 4, 6, 7, 8, 15, 14, 12, 9, 10, 5}. The difference of the values of the first and third quartiles of the data set is

OpenStudy (anonymous):

crap sorry idk

OpenStudy (anonymous):

f#$k u to

OpenStudy (texaschic101):

first put the numbers in order 4,5,6,6,7,8,9,9, 10,10,12,12,14,15 first find quartile 2 (Q2).....it is the median (middle number) if there is an odd number of numbers, then the middle number is easy to find...but if there is an even number, like this one, there will be 2 middle numbers....take the 2 middle numbers and divide by 2.....so Q2 = (9 + 9)/2 = 18/2 = 9 for quartile 1 (Q1)....take all the numbers before Q2 and find the median.. 4,5,6,6,7,8,9.....and take the median (middle number) of that set of numbers. so Q1 = 6 for quartile 3 (Q3)...take all the numbers after Q2 and find the median... 9,10,10,12,12,14,15 so Q3 = 12 so the difference between Q3 and Q1 = 6